# Provable test-time adaptivity and distributional robustness of in-context learning

### v1

@@ Abstract
We study in-context learning problems where a Transformer is pretrained on tasks drawn from a mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$, called the pretraining prior, in which each mixture component $\pi_{\alpha}$ is a distribution on tasks of a specific difficulty level indexed by $\alpha$. Our goal is to understand the performance of the pretrained Transformer when evaluated on a different test distribution $\mu$, consisting of tasks of fixed difficulty $\beta\in\mathcal{A}$, and with potential distribution shift relative to $\pi_{\beta}$, subject to the chi-squared divergence $\chi^{2}(\mu,\pi_{\beta})$ being at most $\kappa$. In particular, we consider nonparametric regression problems with random smoothness, and multi-index models with random smoothness as well as random effective dimension. We prove that a large Transformer pretrained on sufficient data achieves the optimal rate of convergence corresponding to the difficulty level $\beta$, uniformly over test distributions $\mu$ in the chi-squared divergence ball. Thus, the pretrained Transformer is able to achieve faster rates of convergence on easier tasks and is robust to distribution shift at test time. Finally, we prove that even if an estimator had access to the test distribution $\mu$, the convergence rate of its expected risk over $\mu$ could not be faster than that of our pretrained Transformers, thereby providing a more appropriate optimality guarantee than minimax lower bounds.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have emerged as one of the dominant architectures in modern machine learning, as they have achieved state-of-the-art performance in many domains such as natural language processing (Devlin et al., 2019; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; Liu et al., 2021) and protein prediction (Jumper et al., 2021). A striking ability of pretrained Transformers, first observed by Brown et al. (2020) in Large Language Models (LLMs), is the phenomenon of in-context learning (ICL): given a prompt containing examples and a query, Transformers can learn the underlying pattern from the examples and produce accurate output for the query, without updating its parameters. For instance, given the values of the function $(x,y)\mapsto xy^{2}$ or $\text{subfield}\mapsto\text{field}$’ at a few distinct points, LLMs are able to recover the value of the function at a new point:
$\overbrace{\underbrace{f(1,1)=1,\;f(1,2)=4,\;f(2,2)=8,\;f(2,3)=18,}_{\text{examples}}\;\underbrace{f(3,3)=}_{\text{query}}}^{\text{prompt}}\underbrace{27\vphantom{f(2,2)}}_{\text{output}},$ (1)
$\overbrace{\underbrace{\text{Genetics}\to\text{Biology},\;\text{Relativity}\to\text{Physics},}_{\text{examples}}\;\underbrace{\text{Topology}\to}_{\text{query}}}^{\text{prompt}}\underbrace{\text{Mathematics}\vphantom{y}}_{\text{output}}.$
The ICL ability of Transformers was also exploited by Hollmann et al. (2025) to build tabular foundation models, which they show to outperform many classical machine learning methods, such as tree-based and boosting algorithms, on regression and classification problems for tabular data.
Extensive studies on ICL have been conducted from various perspectives (e.g. Garg et al., 2022; Xie et al., 2022; Von Oswald et al., 2023; Bai et al., 2023; Kim, Nakamaki and Suzuki, 2024); we refer the readers to Section 1.1 for a more detailed literature review. However, most theoretical analyses of ICL have focused on the setting where the distribution of pretraining data (which we call the pretraining prior) and the distribution of the test data (which we call the test distribution) are the same. In practice, Transformers are pretrained on a variety of tasks of different difficulties, and the test distribution will typically exhibit distribution shift from the pretraining prior. For example, we may train a Transformer on translation tasks between different languages, where Dutch–English translation is typically easier than Chinese–English, since Dutch and English are both West Germanic languages with shared structure. At test time, the Transformer is required to translate between two specific languages, while some linguistic patterns may appear more frequently in the test data than in the pretraining data. To formulate a theoretical framework, we focus on regression problems (in line with most prior work) and consider a Transformer pretrained on a mixture distribution containing tasks of different difficulties, but where the test distribution consists of tasks of a specific difficulty with potential distribution shift from the corresponding mixture component of the pretraining prior. We study the following natural but important questions:
@@
  • Finally, in Section 4.3, we show that the expected risk of any estimator over the test distribution $\mu$ cannot achieve a faster rate of convergence than our pretrained Transformers (even if the estimator is allowed to depend on $\mu$). This provides a more appropriate form of optimality guarantee for pretrained Transformers than that yielded by minimax lower bounds usually considered in ICL literature.
@@ 1.1 Related work
In-context learning
There is a large body of theoretical research that aims to explain the ICL ability of Transformers. One line of work shows that Transformers can perform ICL by approximating certain algorithms, such as gradient descent (Ahn et al., 2023; Akyürek et al., 2023; Bai et al., 2023; Von Oswald et al., 2023), functional gradient descent (Cheng, Chen and Sra, 2024), basis function regression (Kim, Nakamaki and Suzuki, 2024), reinforcement learning (Lin, Bai and Mei, 2024) and domain adaptation (Hataya, Matsui and Imaizumi, 2024). Other lines of work analyse ICL through the lens of Bayesian inference (Xie et al., 2022; Zhang et al., 2022; Nagler, 2023; Wang et al., 2023; Deora et al., 2025; Zhang et al., 2025), meta-learning (Dai et al., 2023; Jeon et al., 2024; Wu, Wang and Yao, 2025) and training dynamics (Oko et al., 2024; Zhang, Frei and Bartlett, 2024; Kumano, Kera and Yamasaki, 2025; Kuwataka and Suzuki, 2025). In particular, Bai et al. (2023) show that Transformers can perform gradient descent on penalised linear regression, generalised linear models and two-layer neural networks, as well as implement in-context algorithm selection, while Kim, Nakamaki and Suzuki (2024) prove that Transformers are minimax optimal for nonparametric regression. Both papers consider the case where the test distribution coincides with the pretraining prior. On the other hand, Xie et al. (2022) study the setting where the test prompts have different formats compared to the pretraining data, and provide consistency results under a Bayesian hidden Markov model. There are also numerous empirical studies on the behaviour and properties of ICL (e.g. Garg et al., 2022; Kirsch et al., 2022; Panwar, Ahuja and Goyal, 2024; Reuter et al., 2025). Finally, we note the concurrent and independent work of Wakayama and Suzuki (2025), who study the adaptivity of ICL from a Bayesian perspective. Their results are complementary to ours, and neither subsumes the other.
Posterior contraction theory
Posterior contraction theory is an active area of research in statistics that studies the concentration of the posterior distribution around a true data generating distribution (i.e. from a frequentist perspective). Pioneered by Ghosal, Ghosh and van der Vaart (2000) and Shen and Wasserman (2001), posterior contraction for Bayesian nonparametric models has been investigated in different problems (e.g. Ghosal, Lember and van der Vaart, 2008; van der Vaart and van Zanten, 2009; Finocchio and Schmidt-Hieber, 2023; Egels and Castillo, 2025). In this paper, we modify the tools developed by Ghosal, Lember and van der Vaart (2008) to derive rates of convergence for the posterior regression function. Interested readers are referred to the book by Ghosal and van der Vaart (2017) for an introduction to the field.
@@ 1.2 Notation
For $n\in\mathbb{N}$, we write $n]\coloneqq\{1,\ldots,n\}$. For $m,n\in\mathbb{N}$ with $m<n$, we write $m:n]\coloneqq\{m,m+1,\ldots,n\}$ and similarly $n:\infty)\coloneqq\{n,n+1,\ldots\}$. Given $a,b\in\mathbb{R}$, we write $a\wedge b\coloneqq\min\{a,b\}$ and $a\vee b\coloneqq\max\{a,b\}$. For $d\in\mathbb{N}$, we define $\bm{0}_{d}\coloneqq(0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ and $\bm{1}_{d}\coloneqq(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$. For $R>0$, we define the clipping operator $\mathsf{clip}_{R}:\mathbb{R}\to[-R,R]$ by $\mathsf{clip}_{R}(x)\coloneqq-R\vee x\wedge R$ for $x\in\mathbb{R}$. Given a Borel measurable set $A\subseteq\mathbb{R}^{d}$, we write $\mathrm{Vol}_{d}(A)$ for the $d$-dimensional Lebesgue measure of $A$. For two $\sigma$-finite measures $P$ and $Q$ on a measurable space $(\mathcal{X},\Sigma)$, we say that $P$ is absolutely continuous with respect to $Q$, written as $P\ll Q$, if $P(A)=0$ whenever $A\in\Sigma$ is such that $Q(A)=0$, and in that case write $\frac{\mathrm{d}P}{\mathrm{d}Q}$ for the Radon–Nikodym derivative. When $P$ and $Q$ are probability measures, the $\chi^{2}$-divergence from $Q$ to $P$ is defined as
$\chi^{2}(P,Q)\coloneqq\begin{cases}\int_{\mathcal{X}}\bigl(\frac{\mathrm{d}P}{\mathrm{d}Q}\bigr)^{2}\,\mathrm{d}Q-1\quad&\text{if }P\ll Q\\ \infty&\text{otherwise}.\end{cases}$
Further, we write $P^{\otimes n}$ for the $n$-fold product measure of $P$ and $\mathrm{supp}(P)$ for the support of $P$, i.e. the intersection of all closed $A\in\Sigma$ with $P(A)=1$. For a set $\mathcal{F}$ of functions from $\mathcal{X}$ to $\mathcal{Y}$ and a set $\mathcal{G}$ of functions from $\mathcal{Y}$ to $\mathcal{Z}$, we write $\mathcal{F}\circ\mathcal{G}\coloneqq\{f\circ g:f\in\mathcal{F},\,g\in\mathcal{G}\}$, where $f\circ g$ denotes the composition of $f$ and $g$. For a function $f:\mathcal{X}\to\mathbb{R}$, we write $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
@@ 2 Problem set-up and ICL excess risk decomposition
$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{T}(f),\quad\text{where }\widehat{\mathcal{R}}_{T}(f)\coloneqq\frac{1}{T}\sum_{t=1}^{T}\bigl\{Y^{(t)}-f(X^{(t)},\mathcal{D}_{n}^{(t)})\bigr\}^{2}$ (2)
and $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$ are generated according to $\pi$.
We remark that, given a new set of examples $D_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$, the prediction at a query $x\in\mathcal{X}$ is given by $\widehat{f}_{T}(x,D_{n})$, and we do not need to update $\widehat{f}_{T}$ accordingly.
For a distribution $\nu$ on $\mathcal{P}$, we define the $\nu$-risk of a measurable $g:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ as [^1] We adopt the convention that for $P\in\mathcal{P}$, $\mathbb{E}_{P}$ means that the random variables inside the expectation are generated from $P$, i.e. $(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}$, whereas for a distribution $\nu$ on $\mathcal{P}$, $\mathbb{E}_{\nu}\mathbb{E}_{P}$ means that the distribution $P$ is also random and drawn from $\nu$.
$\mathcal{R}_{\nu}(g)$ $\coloneqq\mathbb{E}_{P\sim\nu}\mathbb{E}_{(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]\equiv\mathbb{E}_{\nu}\mathbb{E}_{P}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
@@
$\hskip 2.56064pt=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (3)
where the term $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ is the oracle risk assuming knowledge of the distribution $P$ that generates $(X,Y)$, and in particular of the true regression function $\mathbb{E}_{P}(Y\,|\,X)$ under $P$, and the equality is shown in the proof of Proposition 1. Finally, we define the posterior regression function $g_{\pi}:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ (with respect to $\pi$) by
@@
The following proposition provides a key upper bound on the ICL excess $\mu$-risk, enabling us to study the adaptivity and distributional robustness of ICL.
@@
Let $R>0$, and suppose that for all $P\in\mathcal{P}$, we have $|\mathbb{E}_{P}(Y\,|\,X)|\leq R$ almost surely. Let $\widehat{f}_{T}$ be defined as in (2), let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$ and let $\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})$. Then
@@
The upper bound in (5) consists of two terms. The first involves both the extent of the distribution shift measured by the $\chi^{2}$-divergence, as well as the expected difference between the $\pi$-risk of $\widetilde{f}_{T}$ and that of the posterior regression function $g_{\pi}$. It is well-known that $g_{\pi}$ minimises the $\pi$-risk $\mathcal{R}_{\pi}$ over all measurable functions, so intuitively, if $\widetilde{f}_{T}$ is a large Transformer pretrained on a sufficient corpus of data, then $\mathcal{E}(\widetilde{f}_{T})$ should be negligible. Indeed, this is proved in our Proposition 3. The second term reflects the proximity of the posterior regression function to the true regression function. In fact, in Sections 4.1 and 4.2, we will control $\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ uniformly for all $P$ in the support of $\mu$. This reveals the interesting feature that only the first term on the right-hand side of (5) depends on the degree of distribution shift, and this term can be made arbitrarily small by increasing the size of the Transformer and the pretraining dataset, so that the second term ultimately determines the rate of convergence.

### v2

@@ Abstract
We study in-context learning problems where a Transformer is pretrained on tasks drawn from a mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$, called the pretraining prior, in which each mixture component $\pi_{\alpha}$ is a distribution on tasks of a specific difficulty level indexed by $\alpha$. Our goal is to understand the performance of the pretrained Transformer when evaluated on a different test distribution $\mu$, consisting of tasks of fixed difficulty $\beta\in\mathcal{A}$, and with potential distribution shift relative to $\pi_{\beta}$, subject to the chi-squared divergence $\chi^{2}(\mu,\pi_{\beta})$ being at most $\kappa$. In particular, we consider nonparametric regression problems with random smoothness, and multi-index models with both random smoothness and random effective dimension. We prove that a large Transformer pretrained on sufficient data achieves the optimal rate of convergence corresponding to the difficulty level $\beta$, uniformly over test distributions $\mu$ in the chi-squared divergence ball. Thus, the pretrained Transformer is able to achieve faster rates of convergence on easier tasks and is robust to distribution shift at test time. Finally, we prove that even if an estimator had access to the test distribution $\mu$, the convergence rate of its expected risk over $\mu$ could not be faster than that of our pretrained Transformers, thereby providing a more appropriate optimality guarantee than minimax lower bounds.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have emerged as one of the dominant architectures in modern machine learning, as they have achieved state-of-the-art performance in many domains such as natural language processing (Devlin et al., 2019; Brown et al., 2020), computer vision (Dosovitskiy et al., 2021; Liu et al., 2021) and protein prediction (Jumper et al., 2021). A striking ability of pretrained Transformers, first observed by Brown et al. (2020) in Large Language Models (LLMs), is the phenomenon of in-context learning (ICL): given a prompt (context) containing examples and a query, Transformers can learn the underlying pattern from the examples and produce accurate output for the query, without updating its parameters. The ICL ability of Transformers was also exploited by Hollmann et al. (2025) to build tabular foundation models, which they show to outperform many machine learning methods, such as tree-based and boosting algorithms, on regression and classification problems for tabular data.
GeneticsBiologyRelativityPhysicsTopology?MathematicsEasy taskCat10Fish25Deer17Dog13Bird?9Difficult taskPrompt(Context)Output
Figure 1: Two ICL tasks. Blue indicates the (in-context) examples, red indicates the queries, and green indicates the outputs. The left column is an easy task, while the right column is a harder task which requires more examples to produce accurate output.
[…]
@@ 1.1 Contributions
  • In Section 4.3, we show that the expected risk of any estimator over the test distribution $\mu$ cannot achieve a faster rate of convergence than our pretrained Transformers (even if the estimator is allowed to depend on $\mu$). This provides a more appropriate form of optimality guarantee for pretrained Transformers than that yielded by minimax lower bounds usually considered in ICL literature.
  • Finally, in Section 5, we verify our theoretical results with empirical experiments. These confirm our main conclusions that pretrained Transformers are able to adapt to different task difficulties and are robust to distribution shifts at test time.
@@ 1.2 Related Work
In-context learning   There is a large body of theoretical research that aims to explain the ICL ability of Transformers. One line of work shows that Transformers can perform ICL by approximating certain algorithms, such as gradient descent (Ahn et al., 2023; Akyürek et al., 2023; Bai et al., 2023; von Oswald et al., 2023), functional gradient descent (Cheng, Chen and Sra, 2024), basis function regression (Kim, Nakamaki and Suzuki, 2024), reinforcement learning (Lin, Bai and Mei, 2024) and domain adaptation (Hataya, Matsui and Imaizumi, 2024). Other lines of work analyse ICL through the lens of Bayesian inference (Xie et al., 2022; Zhang et al., 2022; Nagler, 2023; Wang et al., 2023; Deora et al., 2025; Zhang et al., 2025), meta-learning (Dai et al., 2023; Jeon et al., 2024; Wu, Wang and Yao, 2025) and training dynamics (Oko et al., 2024; Zhang, Frei and Bartlett, 2024; Kumano, Kera and Yamasaki, 2025; Kuwataka and Suzuki, 2025). In particular, Bai et al. (2023) show that Transformers can perform gradient descent on penalised linear regression, generalised linear models and two-layer neural networks, as well as implement in-context algorithm selection, while Kim, Nakamaki and Suzuki (2024) prove that Transformers are minimax optimal for nonparametric regression. Both papers consider the case where the test distribution coincides with the pretraining prior. On the other hand, Xie et al. (2022) study the setting where the test prompts have different formats compared to the pretraining data, and provide consistency results under a Bayesian hidden Markov model. There are also numerous empirical studies on the behaviour and properties of ICL (e.g. Garg et al., 2022; Kirsch et al., 2022; Panwar, Ahuja and Goyal, 2024; Reuter et al., 2025). Finally, we note the concurrent and independent work of Wakayama and Suzuki (2025), who study the adaptivity of ICL from a Bayesian perspective. Their results are complementary to ours, and neither subsumes the other.
Posterior contraction theory   Posterior contraction theory is an active area of research in statistics that studies the concentration of the posterior distribution around a true data generating distribution (i.e. from a frequentist perspective). Pioneered by Ghosal, Ghosh and van der Vaart (2000) and Shen and Wasserman (2001), posterior contraction for Bayesian nonparametric models has been investigated in different problems (e.g. Ghosal, Lember and van der Vaart, 2008; van der Vaart and van Zanten, 2009; Finocchio and Schmidt-Hieber, 2023; Egels and Castillo, 2025). In this paper, we modify the tools developed by Ghosal, Lember and van der Vaart (2008) to derive rates of convergence for the posterior regression function. Interested readers are referred to the book by Ghosal and van der Vaart (2017) for an introduction to the field.
@@ 1.3 Notation
For $n\in\mathbb{N}$, we write $n]\coloneqq\{1,\ldots,n\}$. For $m,n\in\mathbb{N}$ with $m<n$, we write $m:n]\coloneqq\{m,m+1,\ldots,n\}$ and similarly $n:\infty)\coloneqq\{n,n+1,\ldots\}$. Given $a,b\in\mathbb{R}$, we write $a\wedge b\coloneqq\min\{a,b\}$ and $a\vee b\coloneqq\max\{a,b\}$. For $d\in\mathbb{N}$, we define $0_{d}\coloneqq(0,\ldots,0)^{\top}\in\mathbb{R}^{d}$, $1_{d}\coloneqq(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$ and $\mathbb{B}^{d}\coloneqq\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$. For $R>0$, we define the clipping operator $\mathsf{clip}_{R}:\mathbb{R}\to[-R,R]$ by $\mathsf{clip}_{R}(x)\coloneqq-R\vee x\wedge R$ for $x\in\mathbb{R}$. Given a Borel measurable set $A\subseteq\mathbb{R}^{d}$, we write $\mathrm{Vol}_{d}(A)$ for the $d$-dimensional Lebesgue measure of $A$. For two $\sigma$-finite measures $P$ and $Q$ on a measurable space $(\mathcal{X},\Sigma)$, we say that $P$ is absolutely continuous with respect to $Q$, written as $P\ll Q$, if $P(A)=0$ for all $A\in\Sigma$ such that $Q(A)=0$, and in that case write $\frac{\mathrm{d}P}{\mathrm{d}Q}$ for the Radon–Nikodym derivative. When $P$ and $Q$ are probability measures, the $\chi^{2}$-divergence from $Q$ to $P$ is defined as $\chi^{2}(P,Q)\coloneqq\int_{\mathcal{X}}\bigl(\frac{\mathrm{d}P}{\mathrm{d}Q}\bigr)^{2}\,\mathrm{d}Q-1$ if $P\ll Q$, and $\chi^{2}(P,Q)\coloneqq\infty$ otherwise. Further, we write $P^{\otimes n}$ for the $n$-fold product measure of $P$ and $\mathrm{supp}(P)$ for the support of $P$, i.e. the intersection of all closed $A\in\Sigma$ with $P(A)=1$. For a set $\mathcal{F}$ of functions from $\mathcal{X}$ to $\mathcal{Y}$ and a set $\mathcal{G}$ of functions from $\mathcal{Y}$ to $\mathcal{Z}$, we write $\mathcal{F}\circ\mathcal{G}\coloneqq\{f\circ g:f\in\mathcal{F},\,g\in\mathcal{G}\}$, where $f\circ g$ denotes the composition of $f$ and $g$. For a function $f:\mathcal{X}\to\mathbb{R}$, we write $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
@@ 2 Problem Set-up and ICL Excess Risk Decomposition
$\widehat{f}_{T}\in\operatorname*{argmin}_{f\in\mathcal{F}}\widehat{\mathcal{R}}_{T}(f),$ (1)
where $\widehat{\mathcal{R}}_{T}(f)\coloneqq\frac{1}{T}\sum_{t=1}^{T}\bigl\{Y^{(t)}-f(X^{(t)},\mathcal{D}_{n}^{(t)})\bigr\}^{2}$ and $(\mathcal{D}_{n}^{(t)},X^{(t)},Y^{(t)})_{t=1}^{T}$ are generated according to $\pi$. We remark that, given a new set of examples $D_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$, the prediction at a query $x\in\mathcal{X}$ is given by $\widehat{f}_{T}(x,D_{n})$, and we do not need to update $\widehat{f}_{T}$ accordingly.
For a distribution $\nu$ on $\mathcal{P}$, we define the $\nu$-risk of a measurable $g:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ as [^1] We adopt the convention that for $P\in\mathcal{P}$, $\mathbb{E}_{P}$ means that the random variables are generated from $P$, i.e. $(\mathcal{D}_{n},X,Y)\sim P^{\otimes(n+1)}$, whereas for a distribution $\nu$ on $\mathcal{P}$, $\mathbb{E}_{\nu}\mathbb{E}_{P}$ means that the distribution $P$ is also random and drawn from $\nu$.
$\mathcal{R}_{\nu}(g)$ $\coloneqq\mathbb{E}_{P\sim\nu}\mathbb{E}_{(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\bigl[\bigl\{Y\!-\!g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr]$
$\hskip 2.27626pt\equiv\mathbb{E}_{\nu}\mathbb{E}_{P}\bigl[\bigl\{Y-g(X,\mathcal{D}_{n})\bigr\}^{2}\bigr].$
@@
$=\mathbb{E}\,\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{\widetilde{f}_{T}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr],$ (2)
where the term $\mathbb{E}_{\mu}\mathbb{E}_{P}\bigl[\bigl\{Y-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ is the oracle risk assuming knowledge of the distribution $P$ that generates $(X,Y)$, and in particular of the true regression function $\mathbb{E}_{P}(Y\,|\,X)$ under $P$; the equality (2) is shown in the proof of Proposition 1, see (17). Finally, we define the posterior regression function $g_{\pi}:\mathcal{X}\times(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{Y}$ (with respect to $\pi$) by
@@
The following proposition provides a key upper bound on the ICL excess $\mu$-risk, which is crucial for our study of the adaptivity and distributional robustness of ICL.
@@
Let $R>0$, and suppose that for all $P\in\mathcal{P}$, we have $|\mathbb{E}_{P}(Y\,|\,X)|\leq R$ almost surely. Further let $\widehat{f}_{T}$ be defined as in (1), let $\widetilde{f}_{T}\coloneqq\mathsf{clip}_{R}\circ\widehat{f}_{T}$ and let $\mathcal{E}(\widetilde{f}_{T})\coloneqq\mathbb{E}\mathcal{R}_{\pi}(\widetilde{f}_{T})-\mathcal{R}_{\pi}(g_{\pi})$. Then we have
@@
The upper bound in (4) consists of two terms. The first involves both the extent of the distribution shift measured by the $\chi^{2}$-divergence, as well as the expected difference between the $\pi$-risk of $\widetilde{f}_{T}$ and that of the posterior regression function $g_{\pi}$. It is well-known that $g_{\pi}$ minimises the $\pi$-risk $\mathcal{R}_{\pi}$ over all measurable functions, so intuitively, if $\widetilde{f}_{T}$ is a large Transformer pretrained on a sufficient corpus of data, then $\mathcal{E}(\widetilde{f}_{T})$ should be negligible. Indeed, this is proved in our Proposition 3. The second term reflects the proximity of the posterior regression function to the oracle regression function. In fact, in Sections 4.1 and 4.2, we will control $\mathbb{E}_{P}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-\mathbb{E}_{P}(Y\,|\,X)\bigr\}^{2}\bigr]$ uniformly for all $P$ in the support of $\mu$. This reveals the interesting feature that only the first term on the right-hand side of (4) depends on the degree of distribution shift, and this term can be made arbitrarily small by increasing the size of the Transformer and the pretraining dataset, so that the second term ultimately determines the rate of convergence.
Bai et al. (2023) showed that Transformers can approximate algorithm selection based on a train-validation split. Their arguments proceeded via a different risk decomposition to that in Proposition 1, where Transformers are regarded as approximators of the oracle regression functions $\mathbb{E}_{P}(Y\,|\,X)$; see also Kim, Nakamaki and Suzuki (2024). By contrast, in Proposition 1 we view Transformers as approximators of the posterior regression function, which has three key advantages. First, we are able to control the ICL excess risk under distribution shift between the pretraining prior and the test distribution. Second, we are able to demonstrate adaptivity of Transformers to the difficulty of the test task, instead of obtaining a rate of convergence that is a weighted average over difficulty levels (typically dominated by the difficult tasks). Finally, by approximating the posterior regression function, Proposition 1 can be applied to settings where no simple closed form algorithms are available or optimal, thereby (in combination with Theorem 2 below) explaining the empirically-observed superiority of Transformer-based ICL compared to classical algorithms for nonparametric regression (e.g. Hollmann et al., 2025).